# LV-MAE: Learning Long Video Representations through Masked-Embedding Autoencoders

### v1

@@ Abstract
In this work, we introduce long-video masked-embedding autoencoders (LV-MAE), a self-supervised learning framework for long video representation. Our approach treats short- and long-span dependencies as two separate tasks. Such decoupling allows for a more intuitive video processing where short-span spatiotemporal primitives are first encoded and are then used to capture long-range dependencies across consecutive video segments. To achieve this, we leverage advanced off-the-shelf multimodal encoders to extract representations from short segments within the long video, followed by pre-training a masked-embedding autoencoder capturing high-level interactions across segments. LV-MAE is highly efficient to train and enables the processing of much longer videos by alleviating the constraint on the number of input frames. Furthermore, unlike existing methods that typically pre-train on short-video datasets, our approach offers self-supervised pre-training using long video samples (e.g., 20+ minutes video clips) at scale. Using LV-MAE representations, we achieve state-of-the-art results on three long-video benchmarks – LVU, COIN, and Breakfast – employing only a simple classification head for either attentive or linear probing. Finally, to assess LV-MAE pre-training and visualize its reconstruction quality, we leverage the video-language aligned space of short video representations to monitor LV-MAE through video-text retrieval.
@@ 1 Introduction
Recently, approaches for handling long-video content have been proposed based on state-space models (SSMs) to address some of the above challenges. ViS4mer [21] utilizes a transformer encoder to extract patch-level features from each individual frame, followed by an efficient aggregation of these representations across all frames using the S4 decoder [14]. VideoMamba [27], a fully SSM-based video model, directly processes spatio-temporal patches from video frames using multiple bidirectional Mamba blocks [55]. These architectures successfully bypass the quadratic complexity of the transformer’s attention mechanism and are trained on sequences of up to 64 frames [21, 27]. Yet their computational demands rise significantly when applied to videos of extended duration.
@@
Our contributions are summarized as follows: (1) We introduce LV-MAE, a method for learning long-range dependencies across short-video segments by pre-training masked-embedding autoencoders on very long video samples (e.g., 20+ minutes). Our method is highly efficient to train and enables the processing of much longer videos by alleviating the constraint on the number of frames. To the best of our knowledge, this is the first approach to apply masked autoencoders on sequences of embedding vectors instead of image patches. (2) Leveraging LV-MAE representations, we achieve state-of-the-art results with minimal fine-tuning, using techniques such as attentive or linear probing, on three long-video downstream tasks: LVU [45], COIN [40], and Breakfast [23]. (3) By exploiting the video-language shared space of the short-video encoder, we introduce an interpretable technique to visualize how semantic knowledge is reconstructed within the LV-MAE encoder.
@@ 2 Related Work
VideoMAE [41] and VideoMAE V2 [42] extend masked autoencoders to video data by reconstructing masked video patches, enabling efficient self-supervised learning. However, they focus on short video clips up to 32 frames, and scaling to longer videos is computationally challenging due to the quadratic complexity of attention mechanisms with respect to sequence length. In contrast to these methods, our work addresses long-video understanding by operating on sequences of embeddings and aim to learn long-range dependencies without being constrained by the number of frames.
@@
Understanding long videos, such as movies or instructional content, is challenging due to the computational cost of processing lengthy sequences and the need to model long-range temporal dependencies. State-space models (SSMs) have been proposed to address some of these challenges. Vis4mer [21] combines a transformer encoder for extracting spatial features with a structured state-space sequence model (S4) [14] decoder to efficiently aggregate representations across frames. VideoMamba [27] is a fully SSM-based video model that processes spatiotemporal patches directly using bidirectional Mamba blocks [55]. While SSMs reduce computational complexity compared to standard transformers, operating directly at the frame level does not fully alleviate computational demands, limiting processing to up to 64 frames [21, 27]. Our work focuses on enhancing long-video representations through a novel approach, specifically designed for long-form content. We evaluate our LV-MAE method on long-form video classification and regression tasks, demonstrating its effectiveness in handling extended temporal sequences. A detailed discussion of long-video language understanding methods that leverage Large Language Models (LLMs) is provided in the appendix, as these approaches, while related, address different aspects of the long-video understanding challenge.
@@ 3.1 Short-video Representation
Given a full video $V$, we first segment it into a set of $N$ consecutive short videos (e.g., five seconds long), represented as $\mathcal{V}=\{\mathbf{v}_{i}\}_{i=1}^{N}$. For each segment, we extract $T$ frames, denoted by $\mathbf{v}_{i}\in\mathbb{R}^{T\times C\times H\times W}$, where $C$, $H$, and $W$ are the number of channels, height, and width of the frames, respectively. Each video segment $\mathbf{v}_{i}$ is then processed by a multimodal model, pre-trained for short video-text alignment, such as LanguageBind [54] or InternVideo2 [43]. This process produces a set of short-video embeddings =$\mathcal{E}=\{\mathbf{e}_{i}\}_{i=1}^{N}$, where $\mathbf{e}_{i}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension.
@@ 3.2 Masked-embedding Autoencoder
The decoder then outputs a set of reconstructed embeddings ^=$\hat{\mathcal{E}}=\{\hat{\mathbf{e}}_{i}\}_{i=1}^{N}$, i.e.,
^$\hat{\mathcal{E}}=\text{Decoder}\left(\mathcal{Z},\mathbf{m},\mathcal{P}\right),$ (2)
@@
To handle videos of varying lengths, we adopt techniques from BERT for managing sequences [8]. Each sequence of short-video segments is limited to a maximum of 256 tokens, with shorter sequences padded using a special [PAD] token. During the self-attention process, we employ an attention mask to prevent these padding tokens from influencing model training. This approach enables our framework to handle videos of arbitrary length while maintaining computational efficiency.
@@ 3.3 Training and Data
Our model is pre-trained on diverse datasets across multiple video domains, enhancing its robustness and adaptability to various tasks. Specifically, we pre-train on over 1,000 long-length movies and TV series, encompassing a wide range of genres, styles, and lengths to ensure comprehensive content diversity. In addition, we incorporate three publicly available datasets. FineVideo [12] includes a vast collection of diverse videos with more than 40,000 videos covering a wide spectrum of categories and domains. With thousands of hours of content, FineVideo represents a rich resource for capturing varied human activities, scenes, and interactions. Additionally, we leverage the train set of MovieClips [45] that contains approximately 7,000 video clips, typically ranging from one to three minutes, curated from various movies. It serves as a diverse collection of cinematic content. Finally, ActivityNet [9], a dataset that focuses on complex human activities that are relevant to daily life, offering a broad range of action categories and scenarios. This diverse collection of video data enables us to train our model on a wide range of video domains, ensuring generalizability across different tasks and environments. Moreover, as our framework operates in a self-supervised manner, the inclusion of additional data is straightforward since we do not require manual annotations or labels for training.
@@
To train our model, we employ a dynamic sampling strategy to handle videos of varying lengths. For each video, we randomly sample a duration in seconds, and this ensures that each batch contains videos of different lengths. This approach allows us to train on entire videos or portions of them, providing our model with exposure to both short and long video segments, which is particularly useful for handling complex temporal dependencies in long-form video understanding tasks. In practice, for simplicity, we set each short video segment to a length of five seconds.
@@
Our pre-training method is highly efficient compared to existing approaches. This efficiency is demonstrated by the reduced number of tokens required during training. Specifically, since we leverage short-video representations extracted by off-the-shelf models, each video segment is represented by a single token. Thus, with each segment lasting five seconds, processing a two-minute video with LV-MAE requires only 24 tokens, compared to the 11,760 tokens required by other frame-level methods (60 frames $\times$ 14 $\times$ 14 patches for a standard 16 $\times$ 16 image grid) [27, 41, 42, 21, 31]. As a result, we can pre-train LV-MAE in only 20 hours on 8 $\times$ A10 GPUs while achieving state-of-the-art results on downstream tasks. Our optimized framework makes pre-training on long videos computationally feasible without compromising performance.
@@
[…]
@@
[…]
@@ 3.4 Interpretable Predictions
Given a reconstructed embeddings output by the decoder $\hat{\mathbf{e}}_{j}$, corresponding to the $j$-th video segment, we retrieve the top-matching captions with the highest response by computing the cosine similarity between $\hat{\mathbf{e}}_{j}$ and each caption’s textual embedding, $\mathbf{e}_{i}^{\ell}$:
$\text{sim}(\hat{\mathbf{e}}_{j},\mathbf{e}_{i}^{\ell})=\frac{\langle\hat{\mathbf{e}}_{j},\mathbf{e}_{i}^{\ell}\rangle}{\|\hat{\mathbf{e}}_{j}\|\|\mathbf{e}_{i}^{\ell}\|}.$ (4)
This way, these top-matching retrieved captions can be inspected to ensure that the semantic meaning of the reconstructed masked embeddings is well captured by the decoder. Visualization examples are provided in section 4.4.
@@ 4 Experiments
We present results obtained by our method and compare them to the latest state-of-the-art (SOTA) long video understanding models such as VideoBERT [39], ViS4mer [21], Turbo [15] and VideoMamba [27]. In particular, we show that LV-MAE outperforms previous SOTA methods on downstream tasks from diverse domains such as movies, activites prediction and proceduarl tasks. Additionally, we perfrom thorough ablation studies to assess the effectiveness of various design choices in our approach. We provide implmentation details in the Appendix.
@@ 4.2 Main Results
We note that previous works fine-tune all model weights for each benchmark, whereas in our approach, only a small number of parameters are trained through attentive probing. Nevertheless, LV-MAE achieves superior performance on most tasks. Furthermore, previous works typically pre-train their models on short-video datasets, such as Kinetics [22] (e.g., ViS4mer [21] and VideoMamba [27]). In contrast, our approach enables pre-training on long videos by leveraging short-video segment representations and does not require labels. This allows the model to capture long-range dependencies and learn robust, long-term temporal representations.
@@ 4.3 Main Properties
In Table 3, we compare LP and AP using either InternVideo2 or LanguageBind embeddings on the LVU benchmark, Breakfast, and COIN datasets. For the LVU benchmark, we report the average score for classification tasks, with full results available in the Appendix. Our findings show that attentive probing consistently outperforms linear probing. This is likely because AP’s additional attention mechanism enables it to capture task-specific context more effectively, whereas LP averages information across the sequence, potentially oversmoothing important details. For instance, on the Breakfast dataset, attentive probing proves especially useful in distinguishing between cooking activities that share similar steps but differ in subtle nuances. When using linear probing, averaging the latent representations may smooth out important distinctions between actions, such as cracking eggs directly into a pan versus stirring them first before cooking. These minor variations in the long video are critical in tasks like identifying scrambled versus fried eggs, where actions overlap but diverge in essential details. Attentive probing’s capacity to retain these nuances across the video sequence makes it better suited for tasks that rely on fine-grained distinctions.
@@ 4.4 Interpretable Predictions Experiment
The ability to understand model predictions is essential to facilitate human understanding of what the model learns. In Sec. 3.4 we propose an interpretability technique for our framework. In this section, we show its effectiveness and verify the reconstructions produced by the decoder. In practice, we used the MovieClips dataset to extract captions. For each short-video segment, $v_{i}$, in the dataset, we generate concise annotations, $a_{i}$, using the Claude Sonnet 3.5 LLM [1]. Next, we sample a long video from the dataset and extract its corresponding embeddings, $\mathcal{E}$, masking $40\%$ of them. Our model is fed with the visible embeddings and tasked with reconstructing all embeddings, resulting in the predicted set, $\hat{\mathcal{E}}$. Each annotation is transformed into a language embedding using LanguageBind, which has the same embedding space as $\mathcal{E}$. Using the cosine similarity measure in Eq. 4, we perform retrieval by selecting the most similar textual embeddings. We conduct this retrieval for both the original embeddings, $\mathcal{E}$, and the predicted embeddings, $\hat{\mathcal{E}}$. We provide a scheme of the process in Figure 4.
@@ 5 Discussion and Limitations
Second, unlike methods operating on image patches—where masked token predictions can be easily visualized—our approach operates on embeddings, which are inherently more abstract and less interpretable. This makes it challenging to directly assess the model’s performance on masked token predictions. To address this, we propose an interpretability strategy that leverages captions or generates them when unavailable, providing a proxy for visualizing and understanding the model’s predictions. This approach aids in interpreting predictions, though it remains an approximation rather than a direct visualization of learned embeddings.
@@ 6 Conclusions
While new methods for video understanding appear rapidly, the majority of existing methods are limited to short video clips and work at the frame level, which restricts their ability to scale to longer videos. In this work we develop a novel self-supervised representation learning model for long video understanding leveraging embeddings of a pre-trained model for shorter videos. The benefit of our approach is threefold: (1) it builds upon the well-established design of the masked autoencoder; (2) it harnesses robust pre-trained models for high-quality embedding extraction; and (3) we seamlessly process varying length video ranging from seconds to potentially hour long efficiently. Our empirical studies show that the proposed method performs better than the state-of-the-art approaches or on par on all long video understanding tasks. Looking ahead, we see substantial potential for extending this framework to additional applications, including generation, retrieval, and beyond, enabling broader utilization in long-form video analysis.

### v2

@@ Abstract
In this work, we introduce long-video masked-embedding autoencoders (LV-MAE), a self-supervised learning framework for long video representation. Our approach treats short- and long-span dependencies as two separate tasks. Such decoupling allows for a more intuitive video processing where short-span spatiotemporal primitives are first encoded and are then used to capture long-range dependencies across consecutive video segments. To achieve this, we leverage advanced off-the-shelf multimodal encoders to extract representations from short segments within the long video, followed by pre-training a masked-embedding autoencoder capturing high-level interactions across segments. LV-MAE is highly efficient to train and enables the processing of much longer videos by alleviating the constraint on the number of input frames. Furthermore, unlike existing methods that typically pre-train on short-video datasets, our approach offers self-supervised pre-training using long video samples (e.g., 20+ minutes video clips) at scale. Using LV-MAE representations, we achieve state-of-the-art results on three long-video benchmarks – LVU, COIN, and Breakfast – employing only a simple classification head for either attentive or linear probing. Finally, to assess LV-MAE pre-training and visualize its reconstruction quality, we leverage the video-language aligned space of short video representations to monitor LV-MAE through video-text retrieval. Code is available at https://github.com/amazon-science/lv-mae.
@@ 1 Introduction
Recently, approaches for handling long-video content have been proposed based on state-space models (SSMs) to address some of the above challenges. ViS4mer [21] utilizes a transformer encoder to extract patch-level features from each individual frame, followed by an efficient aggregation of these representations across all frames using the S4 decoder [14]. VideoMamba [27], a fully SSM-based video model, directly processes spatio-temporal patches from video frames using multiple bidirectional Mamba blocks [56]. These architectures successfully bypass the quadratic complexity of the transformer’s attention mechanism and are trained on sequences of up to 64 frames [21, 27]. Yet, their computational demands rise significantly when applied to videos of extended duration.
@@
Our contributions are summarized as follows: (1) We introduce LV-MAE, a method for learning long-range dependencies across short-video segments by pre-training masked-embedding autoencoders on very long video samples (e.g., 20+ minutes).
Our method is highly efficient to train and enables the processing of much longer videos by alleviating the constraint on the number of frames. To the best of our knowledge, this is the first approach to apply masked autoencoders on sequences of embedding vectors instead of image patches. (2) Leveraging LV-MAE representations, we achieve state-of-the-art results with minimal fine-tuning, using techniques such as attentive or linear probing, on three long-video downstream tasks: LVU [46], COIN [40], and Breakfast [23]. (3) By exploiting the video-language shared space of the short-video encoder, we introduce an interpretable technique to visualize how semantic knowledge is reconstructed within the LV-MAE encoder.
@@ 2 Related Work
VideoMAE [41] and VideoMAE V2 [42] extend masked autoencoders to video data by reconstructing masked video patches, enabling efficient self-supervised learning. However, they focus on short video clips up to 32 frames, and scaling to longer videos is computationally challenging due to the quadratic complexity of attention mechanisms with respect to sequence length. A decoupling strategy is explored by CoSeg [43], yet it also operates at the frame level and therefore, it inherits the same scaling bottleneck. In contrast to these methods, our work addresses long-video understanding by operating on sequences of embeddings and aims to learn long-range dependencies without being constrained by the number of frames. Our novelty lies in the granularity of the low-level representations, leveraging clip-level features to enable efficient training, avoiding costly frame processing.
@@
Understanding long videos, such as movies or instructional content, is challenging due to the computational cost of processing lengthy sequences and the need to model long-range temporal dependencies. State-space models (SSMs) have been proposed to address some of these challenges. Vis4mer [21] combines a transformer encoder for extracting spatial features with a structured state-space sequence model (S4) [14] decoder to efficiently aggregate representations across frames. VideoMamba [27] is a fully SSM-based video model that processes spatiotemporal patches directly using bidirectional Mamba blocks [56]. While SSMs reduce computational complexity compared to standard transformers, operating directly at the frame level does not fully alleviate computational demands, limiting processing to up to 64 frames [21, 27]. Our work focuses on enhancing long-video representations through a novel approach, specifically designed for long-form content. We evaluate our LV-MAE method on long-form video classification and regression tasks, demonstrating its effectiveness in handling extended temporal sequences. A detailed discussion of long-video language understanding methods that leverage Large Language Models (LLMs) is provided in App. 12, as these approaches, while related, address different aspects of the long-video understanding challenge.
@@ 3.1 Short-video Representation
Given a full video $V$, we first segment it into a set of $N$ consecutive short videos (e.g., five seconds long), represented as $\mathcal{V}=\{\mathbf{v}_{i}\}_{i=1}^{N}$. For each segment, we extract $T$ frames, denoted by $\mathbf{v}_{i}\in\mathbb{R}^{T\times C\times H\times W}$, where $C$, $H$, and $W$ are the number of channels, height, and width of the frames, respectively. Each video segment $\mathbf{v}_{i}$ is then processed by a multimodal model, pre-trained for short video-text alignment, such as LanguageBind [55] or InternVideo2 [44]. It produces a set of short-video embeddings $\mathcal{E}=\{\mathbf{e}_{i}\}_{i=1}^{N}$, where $\mathbf{e}_{i}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension.
@@ 3.2 Masked-embedding Autoencoder
The decoder then outputs a set of reconstructed embeddings $\hat{\mathcal{E}}=\{\hat{\mathbf{e}}_{i}\}_{i=1}^{N}$, i.e.,
$\hat{\mathcal{E}}=\text{Decoder}\left(\mathcal{Z},\mathbf{m},\mathcal{P}\right),$ (2)
@@
To handle videos of varying lengths and to maintain computational efficiency, we adopt techniques from BERT [8]. Each sequence of short-video segments is limited to a maximum of 256 tokens, with shorter sequences padded using a special [PAD] token. During the self-attention process, we employ an attention mask to prevent these padding tokens from influencing model training.
@@ 3.3 Training and Data
Our model is pre-trained on diverse datasets across multiple video domains, enhancing its robustness and adaptability to various tasks. Specifically, we pre-train on over 1,000 long-length movies and TV series, encompassing a wide range of genres, styles, and lengths to ensure comprehensive content diversity. In addition, we incorporate three publicly available datasets. FineVideo [12] includes a vast collection of diverse videos with more than 40,000 videos covering a wide spectrum of categories and domains. With thousands of hours of content, FineVideo represents a rich resource for capturing varied human activities, scenes, and interactions. Additionally, we leverage the train set of MovieClips [46] that contains approximately 7,000 video clips, typically ranging from one to three minutes, curated from various movies. It serves as a diverse collection of cinematic content. Finally, ActivityNet [9], a dataset that focuses on complex human activities that are relevant to daily life, offering a broad range of action categories and scenarios. Our diverse video dataset supports training across varied domains, enhancing generalization. Since our framework is self-supervised, adding more data is easy and requires no manual labels.
@@
To train our model, we employ a dynamic sampling strategy to handle videos of varying lengths. For each video, we randomly sample a duration in seconds, and this ensures that each batch contains videos of different lengths. It allows us to train on entire videos or portions of them, providing our model with exposure to both short and long video segments, which is particularly useful for handling complex temporal dependencies in long-form video understanding tasks. In practice, for simplicity, we set each short video segment to a length of five seconds. In App. 9 we explore the impact of segment length.
@@
Our pre-training method is highly efficient compared to existing approaches. This efficiency is demonstrated by the reduced number of tokens required during training. Specifically, since we leverage short-video representations extracted by off-the-shelf models, each video segment is represented by a single token. Thus, with each segment lasting five seconds, processing a two-minute video with LV-MAE requires only 24 tokens, compared to the 11,760 tokens required by other frame-level methods (60 frames $\times$ 14 $\times$ 14 patches for a standard 16 $\times$ 16 image grid) [27, 41, 42, 21, 31]. As a result, we can pre-train LV-MAE in only 20 hours on 8 $\times$ A10 GPUs while achieving state-of-the-art results on downstream tasks. Our optimized framework makes pre-training on long videos computationally feasible without compromising performance. Additional implementation details are provided in App. 7.
@@
$\text{LV-MAE}_{\text{InternVideo2}}$(Ours) 62.62 68.15 57.14 58.15 77.78 39.89 53.66 59.6 0.23 2.68
$\text{LV-MAE}_{\text{LanguageBind}}$(Ours) 77.57 71.57 64.28 58.15 72.84 40.95 58.53 63.4 0.23 2.52
@@
$\text{LV-MAE}_{\text{InternVideo2}}$ (Ours) 93.24 92.72
$\text{LV-MAE}_{\text{LanguageBind}}$ (Ours) 91.55 92.42
@@ 3.4 Interpretable Predictions
Given a reconstructed embeddings output by the decoder $\hat{\mathbf{e}}_{j}$, corresponding to the $j$-th video segment, we retrieve the top-matching captions with the highest response by computing the cosine similarity between $\hat{\mathbf{e}}_{j}$ and each caption’s textual embedding, $\mathbf{e}_{i}^{\ell}$. This way, these top-matched retrieved captions can be inspected to ensure that the semantic meaning of the reconstructed masked embeddings is well captured by the decoder. Visualization examples are provided in section 4.4.
@@ 4 Experiments
We present results obtained by our method and compare them to the latest state-of-the-art (SOTA) long video understanding models such as VideoBERT [39], ViS4mer [21], Turbo [15] and VideoMamba [27]. In particular, we show that LV-MAE outperforms previous SOTA methods on downstream tasks from diverse domains such as movies, activity prediction, and procedural tasks. Additionally, we perform thorough ablation studies to assess the effectiveness of various design choices in our approach. We provide implementation details in App. 7.
@@ 4.2 Main Results
We note that previous works fine-tune all model weights for each benchmark, whereas in our approach, only a small number of parameters are trained through attentive probing. Nevertheless, LV-MAE achieves superior performance on most tasks. Furthermore, previous works typically pre-train their models on short-video datasets, such as Kinetics [22] (e.g., ViS4mer [21] and VideoMamba [27]). Our approach enables label-free pre-training on long videos by leveraging short-video segments, capturing long-range dependencies and robust temporal patterns.
@@ 4.3 Main Properties
In Table 3, we compare LP and AP using either InternVideo2 or LanguageBind embeddings on the LVU benchmark, Breakfast, and COIN datasets. For the LVU benchmark, we report the average score for classification tasks, with full results available in App. 8. Our findings show that attentive probing consistently outperforms linear probing. This is likely because AP’s additional attention mechanism enables it to capture task-specific context more effectively, whereas LP averages information across the sequence, potentially oversmoothing important details. For instance, on the Breakfast dataset, attentive probing proves especially useful in distinguishing between cooking activities that share similar steps but differ in subtle nuances. When using linear probing, averaging the latent representations may smooth out important distinctions between actions, such as cracking eggs directly into a pan versus stirring them first before cooking. These minor variations in the long video are critical in tasks like identifying scrambled versus fried eggs, where actions overlap but diverge in essential details. Attentive probing preserves these fine-grained nuances across the sequence, making it more effective for such distinctions.
@@ 4.4 Interpretable Predictions Experiment
The ability to understand model predictions is essential to facilitate human understanding of what the model learns. In Sec. 3.4 we propose an interpretability technique for our framework. In this section, we show its effectiveness and verify the reconstructions produced by the decoder. In practice, we used the MovieClips dataset to extract captions. For each short-video segment, $v_{i}$, in the dataset, we generate concise annotations, $a_{i}$, using the Claude Sonnet 3.5 LLM [1]. Next, we sample a long video from the dataset and extract its corresponding embeddings, $\mathcal{E}$, masking $40\%$ of them. Our model is fed with the visible embeddings and tasked with reconstructing all embeddings, resulting in the predicted set, $\hat{\mathcal{E}}$. Each annotation is transformed into a language embedding using LanguageBind, which has the same embedding space as $\mathcal{E}$. Using the cosine similarity, we perform retrieval by selecting the most similar textual embeddings. We conduct this retrieval for both the original embeddings, $\mathcal{E}$, and the predicted embeddings, $\hat{\mathcal{E}}$. We provide a scheme of the process in Figure 4. The original model retrieval accuracy (R@5) is 75.34%.
@@ 5 Discussion and Limitations
Second, unlike methods operating on image patches, where masked token predictions can be easily visualized, our approach operates on embeddings, which are inherently more abstract and less interpretable. This makes it challenging to directly assess the model’s performance on masked token predictions. To address this, we propose an interpretability strategy that leverages captions or generates them when unavailable, providing a proxy for visualizing and understanding the model’s predictions. This approach aids in interpreting predictions, though it remains an approximation rather than a direct visualization of learned embeddings.
@@ 6 Conclusions
While new video understanding methods are emerging rapidly, most are confined to short clips and operate at the frame level, limiting their scalability to longer content. In this work, we introduce a novel self-supervised representation learning model for long-video understanding. Our method builds on the well-established masked autoencoder design, utilizes high-quality embeddings from pre-trained short-video models, and efficiently handles videos ranging from a few seconds to potentially several hours. This allows us to capture richer temporal dynamics and semantic structures over extended durations. Empirical results show that our model consistently outperforms or matches state-of-the-art approaches on long-video tasks. Looking ahead, our framework offers strong potential for broader applications such as generation, retrieval, and comprehensive long-form analysis.